# That’s So FETCH: Fashioning Ensemble Techniques for LLM Classification in Civil Legal Intake and Referral

### v1

@@ Abstract
Each year millions of people seek help for their legal problems by calling a legal aid program hotline, walking into a legal aid office, or using a lawyer referral service. The first step to match them to the right help is to identify the legal problem the applicant is experiencing. Misdirection has consequences. Applicants may miss a deadline, experience physical abuse, lose housing or lose custody of children while waiting to connect to the right legal help. We introduce and evaluate the FETCH classifier for legal issue classification and describe two methods for improving accuracy: a hybrid LLM/ML ensemble classification method, and the automatic generation of follow-up questions to enrich the initial problem narrative. We employ a novel data set of 419 real-world queries to a nonprofit lawyer referral service. Ultimately, we show classification accuracy (hits@2) of 97.37% using a mix of inexpensive models, exceeding the performance of the current state-of-the-art GPT-5 model. Our approach shows promise in significantly reducing the cost of guiding users of the legal system to the right resource for their problem while achieving high accuracy.
@@ Introduction
Finding the right lawyer for your legal problem can be hard. Those who know a lawyer may discover that they practice a different kind of law than they need. Legal-aid eligible populations may spend significant time identifying the right program to apply to. Others, over the income limits for legal aid, may spend significant time asking friends or call a lawyer referral service.
Getting legal triage or referral wrong means that the applicant may be prevented from getting legal help that they are otherwise entitled to, with consequences as severe as loss of liberty, inability to get protection from domestic violence, loss of housing, or loss of custody of a child; or as simple as loss of time and money speaking to an attorney who cannot help them with their specific legal problem.
One tool for classifying natural language text is a large language model (LLM). Even current state-of-the-art LLMs do not classify with 100% accuracy. “Ensemble” classification, which combines multiple classification methods to produce a single combined score [7], is a method of improving classification accuracy.
Ensemble classification methods may use multiple models in series, with routing (where a less powerful model routes the text to the model that performs best on that kind of text) [10], or using a voting method where all models generate a score that is combined using an algorithm [9]. This paper introduces the FETCH (Fast Ensemble Tagging and Classification Helper) classifier. FETCH is a multi-label weighted ensemble voting classifier that combines LLMs, keyword matching, and traditional ML.
This work was produced in collaboration with the Virginia Legal Aid Society and the Oregon State Bar. Virginia Legal Society (VLAS) is a nonprofit legal aid program that serves the southern portion of the U.S. state of Virginia and receives approximately 18,000 calls to human intake workers each year. The Oregon State Bar (OSB) is the nonprofit mandatory state-wide bar for the US State of Oregon and fields approximately 100,000 inquiries to its bar referral service each year. For several years, they have used a simple online webform to gather information from applicants. Trained staff then manually match applicants to the appropriate legal referral.
Ultimately, we show that a combination of 3 inexpensive LLM models: GPT-5-nano, gemini-2.5-flash, and mistral-small, together with a simple keyword matching system and the traditional ML Spot API [8] can exceed the performance of the current state-of-the-art GPT-5 model at a significant cost and speed savings.
@@ 1.1 AI for access to justice
The application of AI to the domain of access to justice is well established and is now the topic of a biannual workshop co-located with JURIX and ICAIL [15], as discussed extensively in [21]. See, e.g., applications to legal information tools [6, 26], form-filling and expert systems [20, 19, 25], and dispute resolution [24, 4, 5].
@@ 1.2 Ensemble classification of natural language text
Ensemble classification methods, where multiple models are used in sequence or at the same time, have been used for some time to improve accuracy in classification prior to the widespread popularity of large language models, popularized in [7] . See, e.g., [12], discussing early neural network and decision tree classification. LLMs have been applied to ensemble classification in diverse domains, including psychology [14] and law [10, 13, 17].
@@ 1.3 Legal intake and advice
[21] discussed the application of LLMs to the task of legal intake, focusing specifically on threshold eligibility determinations in the legal aid context. The work in this paper builds on and generalizes that prior work to the broader legal classification task.
@@ 2 The legal problem classification task
Classification is central to each of these tasks. And leaving the classification up to the applicant is risky. Applicants may have an idea of the kind of lawyer they need to help with their problem, but that idea may be wrong. For example, a layperson might be shown all 244 categories and then be asked to decide that their problem getting Social Security Disability benefits falls under the category “Administrative Law” (a relatively obscure term for a layperson). Informally, we observed applicants struggle when faced with distinguishing between abstract legal categories, and this is supported by studies such as [27] (showing that only 29% of people experiencing a legal problem recognize it as such) and [11] (emphasizing that self-assessed legal understanding often diverges from actual understanding).
@@ 3 The FETCH classifier
The FETCH classifier is a Python FastAPI application available as a REST-like API. The API is currently consumed by two separate applications: the Oregon State Bar’s online lawyer referral service, and the Virginia Legal Aid Society’s phone-based intake system.
We employ the same prompt with 3 LLMs, which are each prompted with the taxonomy and the applicant’s problem description to determine if:
@@
We experimented with different levels of detail in the taxonomy for the prompt, but settled on an abbreviated level, with simply the name of each category and its parent category.
The results are then scored by the number of times each node appears in the results of one of the models. The models are manually weighted according to past performance and results are ranked accordingly. Finally, the top 2 nodes are shown to the applicant.
@@ 4 Oregon State Bar’s referral form
The first implementation of the FETCH classifier is in the Oregon State Bar’s online referral system (Figure 1). Applicants enter the information about their legal problem in a standard web form. Follow-up questions continue to use a form metaphor rather than simulated chat. The LLM is instructed to provide questions with radio buttons and checkboxes as much as possible, which are rendered by the frontend application. We hypothesize that the form both reduces effort for the applicants and reduces the “uncanny valley” [23] effect that users experience when talking to an LLM [16].
@@ 5.1 Dataset
In our experiment, we employed a novel dataset of 419 human-annotated queries made to the Oregon State Bar Referral Service over a period of a few weeks in 2025. The dataset contains a total of 31,716 words, with a mean problem description length of 74 words. Approximately 300 of the queries were annotated to one of the 15 top levels of the legal taxonomy only, while the 119 remaining queries were annotated to a specific “terminal node” in the OSB taxonomy. Annotation was performed by legal professionals at the Oregon State Bar during the regular operation of their referral program. The dataset was anonymized by the bar referral service prior to our analysis.
@@
Some examples from the dataset:
@@
  • I Have a small sole proprietor owned business in construction (tubs&counterrs) & cleaning. Im having issues being paid. Im looking to find a lawyer that knows how to… do contracts and lien property filing for my company. Im just starting to get a profit for the first time. Now I can’t seem to be paid. Invoices over due. (Approx 50% of them) causing me to do this….. if i could get a Refural please.
    High-stakes problems, from stalking and domestic violence abuse to emergency housing, are well represented in the dataset: [^2] Unlike in [21], we did not observe any instances of model “censorship” or refusal to classify situations involving violence or abuse, a promising indication of the suitability of LLMs for problems in the legal domain.
  • Hello, Here is the information, I have been served with an eviction notice with numerous months of non payment, for several reason in which i won’t go into detail at this time , until i have met with a lawyer.Iam currently unable to work.It is very important that i find a reputable legal aid lawyer. This is an urgent request in which in hope it will be resolved soon.I have a family emergency which is a life threatening illness to attend to […]
  • I was granted a temporary stalking order on Friday, [date]. A contested hearing is scheduled for [date] and I need an attorney to represent me as a petitioner. I would also like further legal advice on taking legal action against someone who is stalking me. The respondent to my stalking order initially made contact with as my client in a job I was working as social services provider. Months after I had stopped working with her as a housing advocate, she […]
@@ 5.2 Taxonomy
Our study uses the 244 node taxonomy built by the Oregon State Bar. The Oregon State Bar operates a general-purpose bar referral service covering all civil and criminal case types. The Oregon State Bar’s taxonomy is hierarchical with two levels, with the top level grouping the terminal nodes into 15 categories, such as “Administrative Law”, “Real Property”, and “Criminal Law”. Terminal nodes include more specific classification, such as “Family Law > QDROs (Qualified Domestic Relations Orders)”, “Real Property > LLTN – ADA issues/Animals/Accommodations”, “Bankruptcy > Farms”. Importantly, the topics covered by the taxonomy are truly national in scope, representing a superset of alternative taxonomies such as the 135-node Legal Matter Specification Standard (commonly referred to as SALI taxonomy) [18]. We also used the results from the Spot API [8], which use the conceptually smaller civil legal aid-focused LIST Taxonomy. [^3] https://taxonomy.legal/
@@ 5.3 Classifiers
The ensemble method was a simple weighted vote, with weights calculated based on the raw performance of each classifier across the training dataset. The classifier was limited to providing no more than the 2 “best” labels based on the weighted ranking.
@@ 5.4 Evaluation method
6 Prompt iteration
At the initial development stage for the prompt, we experimented with different levels of detail in the prompt, settling on simply including the taxonomy labels without additional explanation.
@@ 7.1 Hits@2 results
Table 1 shows the results as scored using hits@2. Notably, the results show that the ensemble method performs best, exceeding the performance of best-of-class GPT-5 alone and greatly exceeding the average performance of each inexpensive model used in the ensemble classifier. GPT-5-nano’s strong performance at a fraction of the cost of GPT-4.1-mini means that we pivoted to including it in the ensemble classification shortly after its release.
@@ 7.3 Errors
Roofing replacement contract; contractor obligations/next steps Real Property > Construction/Contractors General Litigation , Consumer Law LLM error
Neighbor harassment/stalking (incl. bear spray); safety concerns Real Property > Neighbor Disputes/Nuisance General Litigation > Neighbor Disputes/Nuisance, General Litigation > Personal Injury Synonymous
Dispute over dog ownership/possession with ex-partner; access withheld Consumer Law General Litigation > Animal Law, Family Law > General (Divorce/Separation) Ambiguous
Wedding planner using event photos/video for promotion; cease-and-desist requested Consumer Law General Litigation > General Torts/Privacy, Intellectual Property > Trademark/Copyright Ambiguous
Lost trial for mobile-device while driving; seeks appeal Consumer Law Criminal Law > Post Conviction/ Appeals, Criminal Law > Traffic Offenses Human annotation error
Wants help filing a small-claims case Criminal Law Consumer Law > Small Claims Advice, General Litigation Human annotation error
Car accident; likely uninsured; facing costs; asks about payment plans Debtor/Creditor General Litigation > Personal Injury, General Litigation > Property Damage LLM error
Pandemic unemployment benefits issue; denial/overpayment and appeal options Debtor/Creditor Administrative Law > Unemployment, Labor & Employment > General Human annotation error
Physician Assistant practice without supervising MD/DO; licensing/supervision compliance Labor & Employment Administrative Law > Professional Licensing, Business & Corporate > General (Contracts, Business, Organization) Human annotation error
@@
For the two errors we classified as clear “LLM error,” the LLM response was a form of “general litigation” instead of a more specific category. FETCH can optionally generate follow-up questions, and in four of the error results it did generate apparently relevant questions.
@@ 8 Discussion
We have shown that small-model ensembles substantially improve on the performance of older machine learning and keyword-matching approaches to classifying legal problems, and slightly improve on the performance of frontier LLM models. What’s more, they do so with lower latency and at a significantly lower cost, at least $\tfrac{1}{3}$ of the cost of using a frontier model for the same task.
@@
We did not observe any patterns in the errors in our ensemble classifier’s results. They did not appear to be biased towards one type of legal problem over another. We did not observe a strong pattern of the LLM failing to distinguish two similar categories. Importantly, this tells us that enriching the category descriptions or adding additional clarifying instructions in the prompt is not likely to significantly improve classification performance.
The number of apparent human annotation errors together with the ambiguous classifications suggest that the baseline human performance at the classification task may be comparable to the LLM ensemble classification method.
Although we did not directly measure the impact of asking follow-up questions, we theorize that it would further improve LLM classification accuracy.
@@ 8.3 Limitations and risks
  • We did not extensively test the “no legal problem” condition, particularly when the applicant query shows signs of delusion. We have thoughts on how to sensitively escalate such cases to a human that require further testing.
  • We did not examine the real world benefit to an applicant to having quick and accurate classification of their legal problem and immediate referral without the delay of human classification. For example, if this led to faster resolution of the applicant’s problem.
  • Whenever an LLM is involved in questions of judgment, embedded biases [3] must be considered. We note that the strong agreement between human and LLM in this experiment, as well as the diversity of methods and models in the ensemble, reduces this concern. In addition, in the systems we deployed, applicants always have an opportunity for human review.
@@ 9 Conclusion and future work
The ensemble method shows promise to improve performance on a high-stakes classification task. Importantly, our results show that a weighted combination of inexpensive small language models can outperform the current state-of-the-art model, GPT-5, at approximately $\tfrac{1}{3}$ of the cost.
Although the error rate is relatively small in our study, future work could further narrow the gap. A routing method based on perceived confidence or disagreement among the model responses could supplement the inexpensive LLM models with a call to a best-of-class LLM such as GPT-5. We did not study the impact of asking follow-up questions in this experiment, but that bears future testing as well. An approach that has human annotators manually remove critical information from answers would be labor intensive but a solid way to answer this question.
Finally, we always look for information about human baseline performance when deciding if an LLM’s performance is “good enough”. We look forward to finding a way to gather this information in a future study.

### v2

@@ Abstract
Each year millions of people seek help for their legal problems by calling a legal aid program hotline, walking into a legal aid office, or using a lawyer referral service. The first step to match them to the right help is to identify the legal problem the applicant is experiencing. Misdirected applicants may miss a deadline, experience physical abuse, lose housing or lose custody of children while waiting to connect to the right legal help. We introduce and evaluate the FETCH classifier for legal issue classification and describe two methods for improving accuracy: a hybrid LLM/ML ensemble classification method, and the automatic generation of follow-up questions to enrich the initial problem narrative. We employ a novel dataset of 419 real-world queries to a nonprofit lawyer referral service. We achieve 97.37% accuracy (hits@2) with inexpensive models, slightly above GPT-5, while reducing cost of guiding users to the right legal resource.
@@ Introduction
Finding the right lawyer can be hard. Even those who know a lawyer may find they practice in the wrong area, while low-income applicants struggle to identify the right aid program and others must rely on friends or referral services. Getting legal triage or referral wrong means that the applicant may lose access to timely legal help and risk liberty, physical abuse or injury, loss of housing, or custody of a child.
One tool for classifying natural language text is a large language model (LLM). Even current state-of-the-art LLMs do not classify with 100% accuracy. “Ensemble” classification, which combines multiple classification methods to produce a single score [7], is a method of improving classification accuracy.
This paper introduces the FETCH (Fast Ensemble Tagging and Classification Helper) classifier. FETCH is a multi-label weighted ensemble voting classifier that combines LLMs, keyword matching, and traditional ML.
This work was produced in collaboration with the Virginia Legal Aid Society and the Oregon State Bar. Virginia Legal Society (VLAS) receives approximately 18,000 calls to human intake workers each year. The Oregon State Bar (OSB) uses trained staff to field approximately 100,000 inquiries to its bar referral service each year.
Ultimately, we show that a combination of three inexpensive LLM models: GPT-5-nano, gemini-2.5-flash, and mistral-small, together with keywords and the traditional ML Spot API [8] may meet or exceed the performance of the current state-of-the-art GPT-5 model at a significant cost and speed savings.
@@ 1.1 AI for access to justice
The application of AI to the domain of access to justice is well established and is now the topic of a biannual workshop co-located with JURIX and ICAIL [15], as discussed in [21]. See, e.g., applications to legal information tools [6, 26], form-filling and expert systems [20, 19, 25], and dispute resolution [24, 4, 5].
@@ 1.2 Ensemble classification of natural language text
Ensembles have been used to improve accuracy when classifying text for decades [7, 12], and LLMs have extended ensemble techniques to medicine [14] and law [10, 13, 17].
@@ 1.3 Legal intake and advice
[21] applied LLMs to legal intake, specifically threshold eligibility determinations in the legal aid context. The work in this paper builds on and generalizes that prior work.
@@ 2 The legal problem classification task
Classification is central to each of these tasks. Applicants may have an idea of the kind of lawyer they need to help with their problem, but that idea may be wrong. [^2] For example, a layperson might be asked to decide that their problem getting Social Security Disability benefits falls under the category “Administrative Law” (a relatively obscure term for a layperson). Informally, we observed applicants struggle when faced with distinguishing between abstract legal categories, and this is supported by studies such as [27] (showing that only 29% of people experiencing a legal problem recognize it as such) and [11] (emphasizing that self-assessed legal understanding often diverges from actual understanding).
@@ 3 The FETCH classifier
The FETCH classifier is available as a REST-like API. The API is in use by two separate applications: the Oregon State Bar’s online lawyer referral service, and the Virginia Legal Aid Society’s phone-based intake system.
We employ a single prompt across three LLMs. [^3] https://gist.github.com/nonprofittechy/ef36fb8da928f25c60cd0ecb82a80750 The prompt includes the taxonomy and the applicant’s problem description to determine if:
@@
We experimented with different levels of detail in the taxonomy for the prompt, but settled on an abbreviated level, with simply the name of each category and its parent category. Instructions were adjusted iteratively to balance generating follow-up questions against making a best effort to classify the applicant’s problem.
The results are then scored by the number of times each node appears in the results of one of the models. The models are manually weighted according to performance on a smaller 20-query subset of the dataset and results are ranked accordingly. Finally, the top two nodes are shown to the applicant.
@@ 4 Oregon State Bar’s referral form
The first implementation of the FETCH classifier is in the Oregon State Bar’s online referral system (Figure 1). As shown in Fig. 1, applicants interact with the system through a consistent form metaphor, rather than in the form of a chatbot. We hypothesize that the form both reduces effort for the applicants and reduces the “uncanny valley” [23] effect that users experience when talking to an LLM [16].
@@ 5.1 Dataset
In our experiment, we employed a novel dataset of 419 human-annotated queries made to the Oregon State Bar Referral Service over a period of a few weeks in 2025. The dataset contains a total of 31,716 words, with a mean problem description length of 74 words. Approximately 300 of the queries were annotated to one of the 15 top levels of the legal taxonomy only, while the 119 remaining queries were annotated to a specific “terminal node” in the OSB taxonomy. Annotation was performed by legal professionals at the Oregon State Bar during the regular operation of their referral program. The dataset was anonymized by the bar referral service prior to our analysis. [^4] A small subset with additional deidentification is available publicly at https://gist.github.com/nonprofittechy/2aa8bfc7141569c7fdd7591564d8e270
@@
Examples ranged from short and direct:
@@
    To high-stakes problems, from stalking and domestic violence abuse to emergency housing (details redacted for length and privacy): [^5] Unlike in [21], we did not observe any instances of model “censorship” or refusal to classify situations involving violence or abuse, a promising indication of the suitability of LLMs for problems in the legal domain.
  • I was granted a temporary stalking order on […]. A contested hearing is scheduled for […] and I need an attorney […]. I would also like further legal advice on taking legal action against someone who is stalking me. […] Months after I had stopped working with her as a housing advocate, she […]
@@ 5.2 Taxonomy
Our study uses a 244 node taxonomy maintained by the Oregon State Bar, [^6] For the full taxonomy, see https://gist.github.com/nonprofittechy/c8a85248d3835d77193f88d1ff41eda3 but national in its application. This represents a superset of alternative taxonomies such as the 135-node Legal Matter Specification Standard (commonly referred to as SALI taxonomy) [18] and the smaller civil legal aid-focused LIST Taxonomy. [^7] https://taxonomy.legal/
@@ 5.3 Classifiers
The ensemble method was a simple weighted vote, with weights calculated based on the raw performance of each classifier across a subset of the training dataset. The classifier was limited to providing no more than the 2 “best” labels based on the weighted ranking.
@@ 6.1 Hits@2 results
Table 1 shows the results as scored using hits@2. The keyword classifier set a baseline at 54%, while TF-IDF fared worst. Notably, the results show that the ensemble method performs best, meeting or slightly exceeding the performance of best-of-class GPT-5 alone [^9] While suggestive, the .0.71 margin between GPT-5 and the ensemble is likely to be statistically insignificant. and greatly exceeding the average performance of each inexpensive model used in the ensemble classifier.
@@ 6.3 Errors
Roofing replacement contract; contractor obligations/next steps Realty > Construction General, Consumer LLM error
Neighbor harassment/stalking (incl. bear spray); safety concerns Realty > Neighbor Disputes General > Neighbor Disputes/, General > Personal Injury Synonymous
Dispute over dog ownership/possession with ex-partner; access withheld Consumer General > Animal , Family > Divorce Ambiguous
Wedding planner using event photos/video for promotion; cease-and-desist requested Consumer General > Torts/Privacy, Intellectual Property > Trademark Ambiguous
Lost trial for mobile-device while driving; seeks appeal Consumer Criminal > Appeals, Criminal > Traffic Offenses Human error
Wants help filing a small-claims case Criminal Consumer > Small Claims Advice, General Human error
Car accident; likely uninsured; facing costs; asks about payment plans Debtor/Creditor General > Personal Injury, General > Property Damage LLM error
Pandemic unemployment benefits issue; denial/overpayment and appeal options Debtor/Creditor Administrative > Unemployment, Labor & Employment > General Human error
Physician Assistant practice without supervising MD/DO; licensing/supervision compliance Labor & Employment Administrative > Professional Licensing, Business > General Human error
@@
For the two errors we classified as clear “LLM error,” the LLM response was a form of “general litigation” instead of a more specific category. FETCH can optionally generate follow-up questions, and in four of the error results it did generate apparently relevant questions. We therefore examined how these automatically generated questions perform in practice.
6.4 Follow-up questions for ambiguous queries
Many real-world applicant queries are short, vague, or ambiguous. In other cases, while the applicant may describe a situation with multiple legal issues, they may only be able to find an attorney to help with one. Automatic follow-up questions can help clarify the applicant’s intent and make legal distinctions that the applicant may not be aware of.
6.4.1 Methodology
The FETCH classifier generates up to three follow-up questions when confidence is low. The applicant’s response is then submitted for an enriched attempt at classification.
To qualitatively test the value of this follow-up question mechanism, we selected 3 candidate questions from the dataset and treated them as follows: one question had details artificially removed. A second question was unchanged, but had been classified by OSB as someone without a legal problem. A third question was misclassified by the LLM. Then we reviewed the output of FETCH for relevance and helpfulness of the follow-up questions to reach an appropriate classification. The results are analyzed in Table 4
Query Generated Follow-up Questions Analysis
I need help with my baby. There’s a domestic situation 1. Is this about the baby’s parents or guardians who are or may be divorcing or separating?
2. Is there safety risk now for you or your baby, or are you worried about safety? (Options: Yes, there is immediate danger, There is some risk but not immediate, No risk currently, Not sure)
3. Do you need help with child custody or visitation for your baby? (Options: Yes, No, Not sure)
(Context removed). The generated questions identify the two potential legal issues: divorce and restraining order, and a correct answer would match the original classification.
I have degenerative arthritis in my ankle and have to wear a brace, making it hard to walk unless I use a walker 1. Are you seeking financial assistance or benefits due to your condition?
2. Has your condition affected your ability to work or keep a job?
3. Are you experiencing any problems with accessibility in public places or at your home because of your difficulty walking?
(full question, no attorney assigned by OSB). The clarifying questions are relevant and identify potential legal issues, such as denial of benefits or Americans with Disabilities Act compliance.
There was a car accident […] I saw a letter, explaining I owe \$10,000 due to the car crash, […] if any payment plans could be done and I need a lawyer for legal advice. 1. What is the current status of this case? (Options: Lawsuit filed or court filing, Demand letter or collection letter only, Not sure/unknown)
2. Were you at fault in the car accident? (Options: Yes, No, Unsure)
3. Was there a police report about the accident?
(full question, misclassified by LLM as dealing with the car crash) The clarifying questions do get us closer to the human annotator’s decision that despite originating with a car accident, at core this is a debtor/creditor problem.
Table 4: Examples of ambiguous applicant queries and clarifying follow-up questions generated by FETCH.
@@ 7 Discussion
We have shown that small-model ensembles substantially improve on the performance of older machine learning and keyword-matching approaches to classifying legal problems, and meet or slightly improve on the performance of frontier LLM models. What’s more, they do so with lower latency and at a significantly lower cost, at least $\tfrac{1}{3}$ of the cost of using a frontier model for the same task. There was no evidence of bias in the distribution of errors and safety concerns revealed by the errors appear to be minimal.
@@
We did not observe any patterns in the errors in our ensemble classifier’s results. They did not appear to be biased towards one type of legal problem over another, and similar categories were not consistently confused by the ensemble. Importantly, this tells us that enriching the category descriptions or adding additional clarifying instructions in the prompt is not likely to significantly improve classification performance.
Overall, annotation errors and ambiguities suggest human baseline performance may match the ensemble. Initial qualitative results suggest the follow-up questions would further improve the performance of the system in real-world usage.
@@ 7.3 Limitations and risks
  • While a relatively large dataset in this domain, 419 queries may not adequately cover the 244 nodes in the full taxonomy. Our findings are strongest when considering the 15 top-level categories.
  • During our study, OSB was only able to collect 19 “no legal problem” scenarios, not enough for a robust analysis. Therefore we did not fully test the “no legal problem” condition, especially delusional queries, which will require future sensitivity testing.
  • We did not measure whether faster automated referral improved real outcomes.
  • While bias is important to consider when using LLMs [3], the low error rate, as well as the diversity of methods and models in the ensemble, reduces this concern. Built-in human review is an additional safeguard in this system.
@@ 8 Conclusion and future work
The ensemble method shows promise to improve performance on a high-stakes classification task. Importantly, our results show that a weighted combination of inexpensive small language models can slightly outperform the current state-of-the-art model, GPT-5, at approximately $\tfrac{1}{3}$ of the cost.
Although the error rate is small in our study, future work could further narrow the gap. A routing method based on perceived confidence or disagreement among the model responses could supplement the inexpensive LLM models with a call to a best-of-class LLM such as GPT-5. We additionally hope to run a broader, quantitative study of the value of follow-up answers in improving classification. One approach would be to manually remove information from real queries. Gathering information about the baseline human error rate would add additional value.